Professor’s Annuity Corp. offers a lifetime annuity to retiring professors. For a payment of $90,000 at age 65, the firm will pay the retiring professor $850 a month until death.

a. If the professor’s remaining life expectancy is 15 years, what is the monthly interest rate on this annuity? (Do not round intermediate calculations. Enter your answer as a percent rounded to 2 decimal places.)

b. What is the effective annual interest rate? (Use the monthly rate computed in part (a) rounded to 2 decimal places when expressed as a percent. Enter your answer as a percent rounded to 2 decimal places.)

c. If the monthly interest rate is 0.75%, what monthly annuity payment can the firm offer to the retiring professor? (Do not round intermediate calculations. Round your answer to 2 decimal places.)
Answer #1

1.

Using financial calculator

N=12*15

PMT=850

PV=-90000

FV=0

CPT I/Y=0.65%

2.

=(1+0.65%)^12-1
=8.08%

3.

Using financial calculator

N=12*15

PV=-90000

FV=0

I/Y=0.75%
CPT PMT=912.8399257